# On the topological ranks of Banach -algebras associated with groups of subexponential growth

### v2

@@ 2.1  Topological stable rank
${\rm Lg}_{n}(\mathfrak{B}):=\left\{(b_{1},\ldots,b_{n})\in\mathfrak{B}^{n}\mid\text{the left ideal generated by $b_{1},\ldots,b_{n}$ coincides with $\mathfrak{B}$}\right\}.$
The abbreviation ‘Lg’ stands for ‘left generators’. If $\mathfrak{B}$ has an involution, we also denote ${\rm Lg}_{n}(\mathfrak{B})_{\rm sa}:={\rm Lg}_{n}(\mathfrak{B})\cap\mathfrak{B}^{n}_{\rm sa},$ for the set of self-adjoint left generators. We now give the definition of (left) topological stable rank, as given by Rieffel in [Ri83, Definition 1.4].
@@
Rieffel observed that one gets the same notion if we decide to work with right ideals on a Banach -algebra, at least when the involution is continuous [Ri83, Proposition 1.6]. All the results we will prove here also follow for the right-sided topological stable rank either by virtue of this observation or by analogy. In any case, let us now introduce a characterization for the tuples in ${\rm Lg}_{n}(\mathfrak{B})$ that is easier to work with. It is well known to experts, but we include it here for completeness (and to highlight the fact that it uses $C^{*}$-techniques). In fact, it is sometimes taken as the definition (cf. [Th21, pgs. 16-17]).
@@
If $\sum_{i=1}^{n}b_{i}^{*}b_{i}$ is invertible, then the left ideal generated by $b_{1},\ldots,b_{n}$ contains an invertible element, and so it coincides with $\mathfrak{B}$. On the other hand, and for the sake of contradiction, suppose that $b_{1},\ldots,b_{n}$ generates $\mathfrak{B}$ as a left ideal, but $\sum_{i=1}^{n}b_{i}^{*}b_{i}$ is not invertible.
@@
Now let $\pi:\mathfrak{B}\to\mathbb{B}(\mathcal{H})$ be a faithful -representation. Then
@@
Let ${\rm tsr}(\mathfrak{A})=n<\infty$, $\varepsilon>0$ and $b_{i}\in\mathfrak{B}$, for $1\leq i\leq n$. Then there exist elements $a_{i},a^{\prime}_{i},c_{i}\in\mathfrak{A}$, $1\leq i\leq n$, such that $\lVert a^{\prime}_{i}-b_{i}\rVert_{\mathfrak{B}}<\varepsilon/2$, $\lVert a_{i}-a^{\prime}_{i}\rVert_{\mathfrak{B}}<\varepsilon/2$ and
@@
We now prove the main theorem of the section. The rough idea is to use the characterization of ${\rm Lg}_{n}(\mathfrak{B})$ in terms of invertibility and use Theorem 2.4 to ensure that the inverses stay inside of $\ell^{1}(G\,|\,{\mathscr{C}})$. Lemma 2.3 then provides us with a way of upgrading the convergence in ${\rm C^{*}}(G\,|\,{\mathscr{C}})$ to convergence in $\ell^{1}(G\,|\,{\mathscr{C}})$.
@@
Suppose that $G$ has subexponential growth. Let $\mathfrak{B}$ be a Banach -algebra such that $\ell^{1}(G\,|\,{\mathscr{C}})\subset\mathfrak{B}\subset{\rm C^{*}}(G\,|\,{\mathscr{C}})$. Then ${\rm tsr}(\mathfrak{B})={\rm tsr}\big{(}\ell^{1}(G\,|\,{\mathscr{C}})\big{)}.$
@@
It is enough to prove the unital case, as passing to the unitization still gives a Fell bundle algebra that satisfies $\ell^{1}(G\,|\,\widetilde{\mathscr{C}})\subset\widetilde{\mathfrak{B}}\subset{\rm C^{*}}(G\,|\,\widetilde{\mathscr{C}})$. Applying Lemma 2.7, we have
@@
Indeed, assume that ${\rm tsr}\big{(}{\rm C^{*}}(G\,|\,{\mathscr{C}})\big{)}=n<\infty$, let $\varepsilon>0$ and let $\Phi_{1},\ldots,\Phi_{n}\in C_{\rm c}(G\,|\,\mathscr{C})$ with $M=\max_{1\leq i\leq n}\lVert\Phi_{i}\rVert_{1}$ and use Lemma 2.3 to find a finitely supported positive definite function $\varphi:G\to\mathbb{C}$ such that
@@
Now we use the definition of topological stable rank to find $\Psi_{1},\ldots,\Psi_{n}\in{\rm C^{*}}(G\,|\,{\mathscr{C}})$ that generate ${\rm C^{*}}(G\,|\,{\mathscr{C}})$ as a left ideal and such that
@@
(2.2)${\leq}\lVert\Psi_{i}-\Phi_{i}\rVert_{\infty}|{\rm Supp}(\varphi)|+\frac{\varepsilon}{2}$
@@
<(2.3)${<}\frac{3}{4}\varepsilon<\varepsilon.$
In particular, we have
@@
(2.4)${\leq}(2\max_{1\leq i\leq n}\lVert\Phi_{i}\rVert_{1}+\varepsilon)\sum_{i=1}^{n}\lVert\Psi_{i}-T_{\varphi}\Psi_{i}\rVert_{*},$
@@
So, if $\varepsilon$ is small enough, $\sum_{i=1}^{n}(T_{\varphi}\Psi_{i})^{*}*T_{\varphi}\Psi_{i}$ is close to $\sum_{i=1}^{n}\Psi_{i}^{*}*\Psi_{i}$, so it has to be invertible. But $\sum_{i=1}^{n}(T_{\varphi}\Psi_{i})^{*}*T_{\varphi}\Psi_{i}\in C_{\rm c}(G\,|\,\mathscr{C})$ and, applying by Theorem 2.4, we see that it is invertible in $\ell^{1}(G\,|\,\mathscr{C})$. Reasoning as before, we conclude that $T_{\varphi}\Psi_{1},\ldots,T_{\varphi}\Psi_{n}$ generates $\ell^{1}(G\,|\,\mathscr{C})$ as a left ideal. Hence $(\Phi_{1},\ldots,\Phi_{n})$ belongs to the closure of ${\rm Lg}_{n}\big{(}\ell^{1}(G\,|\,\mathscr{C})\big{)}$ and this proves that ${\rm tsr}\big{(}\ell^{1}(G\,|\,\mathscr{C})\big{)}\leq n$. ∎
@@
Finally, in the case that $\mathfrak{B}$ is only a Banach algebra, we can only guarantee the equality in the case of topological stable rank 1.
@@
Suppose that $G$ has subexponential growth and that ${\rm tsr}\big{(}{\rm C^{*}}(G\,|\,{\mathscr{C}})\big{)}=1$. Let $\ell^{1}(G\,|\,{\mathscr{C}})\subset\mathfrak{B}$ be any continuous, dense inclusion into a Banach algebra $\mathfrak{B}$. If $\ell^{1}(G\,|\,{\mathscr{C}})$ is unital, we also assume that the inclusion $\ell^{1}(G\,|\,{\mathscr{C}})\subset\mathfrak{B}$ is unital. Then ${\rm tsr}(\mathfrak{B})=1.$
@@
${\rm tsr}(\mathfrak{B})\leq{\rm tsr}\big{(}\ell^{1}(G\,|\,{\mathscr{C}})\big{)}={\rm tsr}\big{(}{\rm C^{*}}(G\,|\,{\mathscr{C}})\big{)}=1,$
@@ 2.2  Real rank
We now proceed to our study of the real rank. The reader will notice that the similarity of definitions allows our line of argumentation to stay essentially the same, provided that we take care of self-adjointness and the corresponding shift in indices.
@@
The proof is the same as the proof of Lemma 2.7, but we include it here. Let ${\rm rr}(\mathfrak{A})=n<\infty$, $\varepsilon>0$ and $b_{i}\in\mathfrak{B}_{\rm sa}$, for $0\leq i\leq n$. Then there exists elements $a_{i},a^{\prime}_{i}\in\mathfrak{A}_{\rm sa}$, $c_{i}\in\mathfrak{A}$, with $0\leq i\leq n$, such that $\lVert a^{\prime}_{i}-b_{i}\rVert_{\mathfrak{B}}<\varepsilon/2$, $\lVert a_{i}-a^{\prime}_{i}\rVert_{\mathfrak{B}}<\varepsilon/2$ and
@@
Suppose that $G$ has subexponential growth. Let $\mathfrak{B}$ be a Banach -algebra with continuous involution and such that $\ell^{1}(G\,|\,{\mathscr{C}})\subset\mathfrak{B}\subset{\rm C^{*}}(G\,|\,{\mathscr{C}})$. Then ${\rm rr}(\mathfrak{B})={\rm rr}\big{(}\ell^{1}(G\,|\,{\mathscr{C}})\big{)}.$
@@
As before, it is enough to prove the unital case and, because of Lemma 2.14, we only need to prove that ${\rm rr}\big{(}{\rm C^{*}}(G\,|\,{\mathscr{C}})\big{)}\geq{\rm rr}\big{(}\ell^{1}(G\,|\,{\mathscr{C}})\big{)}$.
Suppose that ${\rm rr}\big{(}{\rm C^{*}}(G\,|\,{\mathscr{C}})\big{)}=n<\infty$. We let $\varepsilon>0$ and $\Phi_{0},\ldots,\Phi_{n}\in C_{\rm c}(G\,|\,\mathscr{C})_{\rm sa}$ with $M=\max_{0\leq i\leq n}\lVert\Phi_{i}\rVert_{1}$ and use Lemma 2.3 to find a finitely supported positive definite function $\varphi:G\to\mathbb{C}$ such that
@@
For small $\varepsilon$, $\sum_{i=0}^{n}(T_{\varphi}\Psi_{i})^{2}$ will be close to $\sum_{i=0}^{n}\Psi_{i}^{2}$, so it will have to be invertible. But $\sum_{i=0}^{n}(T_{\varphi}\Psi_{i})^{2}\in C_{\rm c}(G\,|\,\mathscr{C})$, so, again by Theorem 2.4, the inverse lies in $\ell^{1}(G\,|\,\mathscr{C})$. Because $\varphi$ is positive definite, this time $(T_{\varphi}\Psi_{0},\ldots,T_{\varphi}\Psi_{n})\in{\rm Lg}_{n+1}\big{(}\ell^{1}(G\,|\,\mathscr{C})\big{)}_{\rm sa}$ and thus ${\rm rr}\big{(}\ell^{1}(G\,|\,\mathscr{C})\big{)}\leq n$. ∎
@@ 3 Applications to symmetrized twisted $L^{p}$-crossed products
As mentioned before, the purpose of this section is to apply the previous results to the case of symmetrized $L^{p}$-crossed products. For that purpose, we will fix a twisted $C^{*}$-dynamical system $(G,C_{0}(X),\alpha,\omega)$, where $G$ is a group of subexponential growth, $X$ a Hausdorff locally compact space, $\alpha$ an action of $G$ on $X$ and $\omega:G\times G\to\mathcal{U}(C_{b}(X))$ a $2$-cocycle. Here, we abuse the notation and call $\alpha$ both the action on $X$ and the induced action on $C_{0}(X)$. That is
@@
It is well known that if $\mathfrak{B}$ has a contractive approximate unit and $\pi:\mathfrak{B}\to\mathbb{B}({\mathscr{E}})$ is a non-degenerate, contractive representation, then $\pi$ extends uniquely to a (unital) representation $\pi:\mathcal{M}(\mathfrak{B})\to\mathbb{B}({\mathscr{E}})$ of the multiplier algebra of $B$. This is always the case for $C_{0}(X)$.
@@
Let $p\in[1,\infty],$ and let $(G,C_{0}(X),\alpha,\omega)$ be a twisted $C^{*}$-dynamical system as above. Let $(Y,\mathcal{B},\mu)$ be a measure space. A covariant representation of $(G,C_{0}(X),\alpha,\omega)$ on $L^{p}(Y,\mu)$ is a pair $(v,\pi)$ consisting of a map $g\mapsto v_{g}$ from $G$ to the invertible operators on $L^{p}(Y,\mu)$ and a non-degenerate, contractive representation $\pi:C_{0}(X)\to\mathbb{B}(L^{p}(Y,\mu)),$ such that
@@
$(\pi\rtimes v)(\Phi)(\xi)=\sum_{x\in G}\pi(\Phi(x))(v_{x}(\xi)),$
for all $\Phi\in\ell^{1}_{\alpha,\omega}(G,C_{0}(X))$ and all $\xi\in L^{p}(Y,\mu)$.
It is well known that every representation of $C_{0}(X)$ induces a covariant representation of $(G,C_{0}(X),\alpha,\omega)$. We proceed to describe this process. Indeed, let $\pi_{0}:C_{0}(X)\to\mathbb{B}(L^{p}(Y,\mu))$ be a contractive representation and let $\nu$ denote the counting measure on $G$. Then the following formulas define representations $v$ of $G$ and $\pi$ of $C_{0}(X)$ on $\ell^{p}(G,\,L^{p}(Y,\mu))\cong L^{p}(G\times Y,\nu\times\mu)$:
$v_{y}(\xi)(x)=\pi_{0}(\alpha_{x}^{-1}(\omega(y,y^{-1}x)))\xi(y^{-1}x),\quad\text{ and }\quad\big{(}\pi(f)\xi\big{)}(x)=\pi_{0}(\alpha_{x}^{-1}(f))\big{(}\xi(x)\big{)},$
for all $x,y\in G$, $\xi\in\ell^{p}(G,\,L^{p}(Y,\mu))$, and $f\in C_{0}(X)$.
It is easy to see that the pair $(v,\pi)$ constructed above is a covariant representation, where $v$ is isometric and $\pi$ contractive. It is usually called the regular covariant representation associated with $\pi_{0}$ [CP13, Definition 2.12]. If $\pi_{0}$ is the point-evaluation representation
$\pi_{0}:C_{0}(X)\to\mathbb{C},\quad\text{ given by }\pi_{0}(f)=f(z),$
then we use the notation $(v,\pi_{z})$ for the associated regular covariant representation. In that case, the integrated representation acquires the following form:
$(\pi_{z}\rtimes v)(\Phi)\xi(x)=\sum_{y\in G}\Phi(y)(xz)\xi(y^{-1}x)\omega(y,y^{-1}x)(xz),$
for all $x\in G,\Phi\in\ell^{1}_{\alpha,\omega}(G,C_{0}(X))$ and all $\xi\in\ell^{p}(G)$.
With these definitions at hand, we can introduce the following norms on $\ell^{1}_{\alpha,\omega}(G,C_{0}(X))$.
$\|\Phi\|_{p,\lambda}$ $=\sup\{\|(\pi\rtimes v)(\Phi)\|\mid(v,\pi)\text{ is a regular covariant pair}\},$
$\|\Phi\|_{p,r}$ $=\sup\{\|(\pi_{z}\rtimes v)(\Phi)\|\mid(v,\pi_{z})\text{ is a regular covariant pair and }z\in X\}.$
The completion $F^{p}(G,X,\alpha,\omega):=\overline{\ell^{1}_{\alpha,\omega}(G,C_{0}(X))}^{\|\cdot\|_{p,\lambda}}$ is usually called the reduced $L^{p}$-crossed product associated with $(G,C_{0}(X),\alpha,\omega)$. It is easily seen to be a Banach algebra. There is also a ‘full’ crossed product construction, which is obtained using all representations into $L^{p}$-spaces and not just the regular ones, but in our setting (the groups are amenable), these two constructions are known to coincide [BaKw25, Theorem 3.5], so we avoid introducing it.
On the other hand, the norms $\|\cdot\|_{p,r}$ can and have been used to introduce symmetrized versions of the $L^{p}$-crossed product (cf. [AuOr22]). Indeed, we let $p,q\in[1,\infty]$ be Hölder duals and define $F^{q}_{*}(G,X,\alpha,\omega)$ to be the completion of $\ell^{1}_{\alpha,\omega}(G,C_{0}(X))$ under the norm
$\lVert\Phi\rVert_{q,*}=\max\{\|\Phi\|_{p,r},\|\Phi\|_{q,r}\}.$
The point is that $F^{q}_{*}(G,X,\alpha,\omega)$ is naturally a Banach -algebra that sits between $\ell^{1}_{\alpha,\omega}(G,C_{0}(X))$ and $C_{0}(X)\rtimes_{\alpha,\omega}G$, so it fits our setting perfectly. Indeed, the fact that this algebra has a continuous involution comes from the fact that, for all $\Phi\in\ell^{1}_{\alpha,\omega}(G,C_{0}(X)),\xi\in\ell^{p}(G),\eta\in\ell^{q}(G)$, one has
$\langle(\pi_{z}\rtimes v)(\Phi)\xi,\eta\rangle=\langle\xi,(\pi_{z}\rtimes v)(\Phi^{*})\eta\rangle,$

### v3

@@ 2.1 Topological stable rank
[…]
The abbreviation ‘Lg’ stands for ‘left generators’. If $\mathfrak{B}$ has an involution, we also use ${\rm Lg}_{n}(\mathfrak{B})_{\rm sa}:={\rm Lg}_{n}(\mathfrak{B})\cap\mathfrak{B}^{n}_{\rm sa},$ for the set of self-adjoint left generators. We now give the definition of (left) topological stable rank, as given by Rieffel in [33, Definition 1.4].
@@
Rieffel observed that we get the same notion if we decide to work with right ideals on a Banach -algebra, at least when the involution is continuous [33, Proposition 1.6]. All the results we will prove here also hold for the right-sided topological stable rank, either by virtue of this observation or by analogy. In any case, let us now introduce a characterization for the tuples in ${\rm Lg}_{n}(\mathfrak{B})$ that is easier to work with. It is well known to experts, but we include it here for completeness (and to highlight the fact that it uses $C^{*}$-algebraic techniques). In fact, it is sometimes taken as the definition of ${\rm Lg}_{n}(\mathfrak{B})$ (cf. [38, pgs. 16-17]).
@@
If $\sum_{i=1}^{n}b_{i}^{*}b_{i}$ is invertible, then the left ideal generated by $b_{1},\ldots,b_{n}$ contains an invertible element, and so it coincides with $\mathfrak{B}$. On the other hand, and for the sake of contradiction, suppose that $b_{1},\ldots,b_{n}$ generates $\mathfrak{B}$ as a left ideal, but that $\sum_{i=1}^{n}b_{i}^{*}b_{i}$ is not invertible.
@@
Now let $\pi:\mathfrak{B}\to\mathbb{B}(\mathcal{H})$ be a unital faithful -representation. Then
@@
Let ${\rm tsr}(\mathfrak{A})=n<\infty$, let $\varepsilon>0$, and let $b_{i}\in\mathfrak{B}$, for $1\leq i\leq n$. Because the inclusion of $\mathfrak{A}$ in $\mathfrak{B}$ is dense, we can find elements $a^{\prime}_{i}\in\mathfrak{A}$ and $1\leq i\leq n$ such that $\lVert a^{\prime}_{i}-b_{i}\rVert_{\mathfrak{B}}<\varepsilon/2$. Now use the definition of ${\rm tsr}(\mathfrak{A})=n$ and the continuity of $\mathfrak{A}\subset\mathfrak{B}$ to choose elements $a_{i},c_{i}\in\mathfrak{A}$, $1\leq i\leq n$, such that $\lVert a_{i}-a^{\prime}_{i}\rVert_{\mathfrak{B}}<\varepsilon/2$ and
@@
We now prove the main theorem of the section. The underlying idea is to use the characterization of ${\rm Lg}_{n}(\mathfrak{B})$ in terms of invertibility and then appeal to Theorem 2.4 to ensure that the corresponding inverses remain inside $\ell^{1}(G\,|\,{\mathscr{C}})$. Lemma 2.3 then provides us with a way of upgrading the convergence in ${\rm C^{*}}(G\,|\,{\mathscr{C}})$ to convergence in $\ell^{1}(G\,|\,{\mathscr{C}})$.
@@
Suppose that $G$ has subexponential growth. Let $\mathfrak{B}$ be a Banach algebra such that $\ell^{1}(G\,|\,{\mathscr{C}})\subset\mathfrak{B}\subset{\rm C^{*}}(G\,|\,{\mathscr{C}})$. Then ${\rm tsr}(\mathfrak{B})={\rm tsr}\big(\ell^{1}(G\,|\,{\mathscr{C}})\big).$
@@
[…]
@@
Indeed, assume that ${\rm tsr}\big({\rm C^{*}}(G\,|\,{\mathscr{C}})\big)=n<\infty$, let $\varepsilon>0$, and let $\Phi_{1},\ldots,\Phi_{n}\in C_{\rm c}(G\,|\,\mathscr{C})$ with $M=\max_{1\leq i\leq n}\lVert\Phi_{i}\rVert_{1}$. Appealing to Lemma 2.3 and the subsequent discussion regarding the net $(T_{{\varphi_{j}}})_{j}$, we can find a finitely supported positive definite function $\varphi:G\to\mathbb{C}$ such that
@@
Now, we use the definition of topological stable rank to find $\Psi_{1},\ldots,\Psi_{n}\in{\rm C^{*}}(G\,|\,{\mathscr{C}})$ that generates ${\rm C^{*}}(G\,|\,{\mathscr{C}})$ as a left ideal such that
@@
[…]
@@
[…]
In particular, using the inequality in (2.3) again, we see that
@@
[…]
@@
So, if $\varepsilon$ is small enough, $\sum_{i=1}^{n}(T_{\varphi}\Psi_{i})^{*}*T_{\varphi}\Psi_{i}$ is close to $\sum_{i=1}^{n}\Psi_{i}^{*}*\Psi_{i}$, so it has to be invertible. But $\sum_{i=1}^{n}(T_{\varphi}\Psi_{i})^{*}*T_{\varphi}\Psi_{i}\in C_{\rm c}(G\,|\,\mathscr{C})$ and, by applying Theorem 2.4, we see that it is invertible in $\ell^{1}(G\,|\,\mathscr{C})$. Reasoning as before, we conclude that $T_{\varphi}\Psi_{1},\ldots,T_{\varphi}\Psi_{n}$ generates $\ell^{1}(G\,|\,\mathscr{C})$ as a left ideal. Hence, $(\Phi_{1},\ldots,\Phi_{n})$ belongs to the closure of ${\rm Lg}_{n}\big(\ell^{1}(G\,|\,\mathscr{C})\big)$, and this proves that ${\rm tsr}\big(\ell^{1}(G\,|\,\mathscr{C})\big)\leq n$. ∎
@@
Finally, in the case that $\mathfrak{B}$ lacks an involution and is only assumed to be a Banach algebra, we can achieve the desired equality of stable ranks only in the case of topological stable rank 1.
@@
Suppose that $G$ has subexponential growth and let $\ell^{1}(G\,|\,{\mathscr{C}})\subset\mathfrak{B}$ be any continuous, dense inclusion into a Banach algebra $\mathfrak{B}$. Then
${\rm tsr}(\mathfrak{B})\leq{\rm tsr}\big({\rm C^{*}}(G\,|\,{\mathscr{C}})\big).$
In particular, if ${\rm tsr}\big({\rm C^{*}}(G\,|\,{\mathscr{C}})\big)=1$, then ${\rm tsr}(\mathfrak{B})=1.$
@@
${\rm tsr}(\mathfrak{B})\leq{\rm tsr}\big(\ell^{1}(G\,|\,{\mathscr{C}})\big)={\rm tsr}\big({\rm C^{*}}(G\,|\,{\mathscr{C}})\big),$
@@ 2.2 Real rank
We now proceed to our study of the real rank. The reader will notice that the similarity of definitions allows our line of argumentation to remain essentially the same, provided that we take care of self-adjointness and the corresponding shift in indices.
@@
The proof is the same as the proof of Lemma 2.7, but we include it here. Let ${\rm rr}(\mathfrak{A})=n<\infty$, let $\varepsilon>0$, and let $b_{i}\in\mathfrak{B}_{\rm sa}$ for $0\leq i\leq n$.
Because the inclusion of $\mathfrak{A}$ in $\mathfrak{B}$ is dense, we can find self-adjoint elements $a^{\prime}_{i}\in\mathfrak{A}$, $0\leq i\leq n$, such that $\lVert a^{\prime}_{i}-b_{i}\rVert_{\mathfrak{B}}<\varepsilon/2$ (if needed, replace $a^{\prime}_{i}$ with $\tfrac{1}{2}a^{\prime}_{i}+\tfrac{1}{2}(a^{\prime}_{i})^{*}$). Now, we proceed as before and use the definition of ${\rm rr}(\mathfrak{A})=n$ and the continuity of $\mathfrak{A}\subset\mathfrak{B}$ to choose elements $a_{i},c_{i}\in\mathfrak{A}_{\rm sa}$, $0\leq i\leq n$, such that $\lVert a_{i}-a^{\prime}_{i}\rVert_{\mathfrak{B}}<\varepsilon/2$ and
@@
Suppose that $G$ has subexponential growth. Let $\mathfrak{B}$ be a Banach -algebra with a continuous involution and such that $\ell^{1}(G\,|\,{\mathscr{C}})\subset\mathfrak{B}\subset{\rm C^{*}}(G\,|\,{\mathscr{C}})$. Then ${\rm rr}(\mathfrak{B})={\rm rr}\big(\ell^{1}(G\,|\,{\mathscr{C}})\big).$
@@
As before, it is enough to prove the unital case, and because of Lemma 2.14, we only need to prove that ${\rm rr}\big({\rm C^{*}}(G\,|\,{\mathscr{C}})\big)\geq{\rm rr}\big(\ell^{1}(G\,|\,{\mathscr{C}})\big)$.
Suppose that ${\rm rr}\big({\rm C^{*}}(G\,|\,{\mathscr{C}})\big)=n<\infty$. We let $\varepsilon>0$ and $\Phi_{0},\ldots,\Phi_{n}\in C_{\rm c}(G\,|\,\mathscr{C})_{\rm sa}$ with $M=\max_{0\leq i\leq n}\lVert\Phi_{i}\rVert_{1}$. Appealing to Lemma 2.3 and the subsequent discussion regarding the net $(T_{{\varphi_{j}}})_{j}$, we can find a finitely supported positive definite function $\varphi:G\to\mathbb{C}$ such that
@@
For small $\varepsilon$, $\sum_{i=0}^{n}(T_{\varphi}\Psi_{i})^{2}$ will be close to $\sum_{i=0}^{n}\Psi_{i}^{2}$, so it will need to be invertible. But $\sum_{i=0}^{n}(T_{\varphi}\Psi_{i})^{2}\in C_{\rm c}(G\,|\,\mathscr{C})$, again by Theorem 2.4, the inverse lies in $\ell^{1}(G\,|\,\mathscr{C})$. Because $\varphi$ is positive definite, this implies $(T_{\varphi}\Psi_{0},\ldots,T_{\varphi}\Psi_{n})\in{\rm Lg}_{n+1}\big(\ell^{1}(G\,|\,\mathscr{C})\big)_{\rm sa}$, and thus ${\rm rr}\big(\ell^{1}(G\,|\,\mathscr{C})\big)\leq n$. ∎
@@ 3 Applications to symmetrized twisted $L^{p}$-crossed products
As mentioned before, the purpose of this section is to apply the previous results to the case of symmetrized $L^{p}$-crossed products. For that purpose, we will fix a twisted $C^{*}$-dynamical system $(G,C_{0}(X),\alpha,\omega)$, where $G$ is a (discrete) group of subexponential growth, $X$ a Hausdorff locally compact space, $\alpha$ an action of $G$ on $X$, and $\omega:G\times G\to\mathcal{U}(C_{b}(X))$ a $2$-cocycle. Here, we abuse the notation and call $\alpha$ both the action on $X$ and the induced action on $C_{0}(X)$. That is
@@
It is well known that if $\mathfrak{B}$ has a contractive approximate unit and $\pi:\mathfrak{B}\to\mathbb{B}({\mathscr{E}})$ is a non-degenerate, contractive representation on a Banach space $\mathscr{E}$}, then $\pi$ extends uniquely to a (unital) representation $\pi:\mathcal{M}(\mathfrak{B})\to\mathbb{B}({\mathscr{E}})$ of the multiplier algebra of $\mathfrak{B}$. This is always the case for $\mathfrak{B}=C_{0}(X)$, and it is relevant since our twisting $2$-cocycles are $\mathcal{M}(C_{0}(X))=C_{b}(X)$-valued. Note that the representation $\pi:\mathcal{M}(\mathfrak{B})\to\mathbb{B}({\mathscr{E}})$ remains contractive [8, Theorem 3.4].
@@
Let $p\in[1,\infty],$ and let $(G,C_{0}(X),\alpha,\omega)$ be a twisted $C^{*}$-dynamical system as described above. Let $(Y,\mathcal{B},\mu)$ be a measure space. A covariant representation of $(G,C_{0}(X),\alpha,\omega)$ on $L^{p}(Y,\mu)$ is a pair $(v,\pi)$ consisting of a map $g\mapsto v_{g}$ from $G$ to the invertible isometries on $L^{p}(Y,\mu)$ and a non-degenerate, contractive representation $\pi:C_{0}(X)\to\mathbb{B}(L^{p}(Y,\mu)),$ such that
@@
$(\pi\rtimes v)(\Phi)=\sum_{x\in G}\pi(\Phi(x))v_{x},$
for all $\Phi\in\ell^{1}_{\alpha,\omega}(G,C_{0}(X))$.
With these definitions at hand, we can introduce the following norm on $\ell^{1}_{\alpha,\omega}(G,C_{0}(X))$: